A certain violet light has a wavelength of 410 nanometers.
What is the energy content of one quantum of the light?
1
Expert's answer
2020-11-09T13:57:22-0500

"E=hv"

v=4.1×10-9m

h= 6.626×10-34

E=(6.626×10-34).(4.1×10-9)

E=2.72×10-42 J